I have several questions like the following that I cannot seem to get the correct answer.
Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b.
(3,6); x+7y=3.
I am missing a step when I try to solve so if I could get a detailed way to solve this that would be great.
Thank you.

Answer by mananth(16949)

1 x + 7 y = 3
Find the slope of this line
make y the subject
7 y = -1 x + 3
Divide by 7
y = - 1/ 7 x + 3/7
Compare this equation with y=mx+b
slope m = - 1/7
The slope of a line parallel to the above line will be the same
The slope of the required line will be - 1/7
m= - 1/7 ,point ( 3 , 6 )
Find b by plugging the values of m & the point in
y=mx+b
6 = -3/7 + b
b= 45/7
m= - 1/7
Plug value of the slope and b in y = mx +b
The required equation is y = - 1/7 x + 45/7

m.ananth@hotmail.ca